Inequality equation: There are 3 fathers (Mr. Sawyer, Mr. Smith, Mr. Rein) whose ages are no less than 70 (combined). Mr. Smith is 20 years younger than Mr. Sawyer. Mr. Rein is 1/3 the age of Sawyer. What are their ages.
Thank you.

Answer by richwmiller(17219)

r+h+n>=70,
h=r-20,
n=r/3
possible answers
h=16, 19, 22,25,28, 31, 34 add 3 to last (34)
r=36, 39, 42,45,48, 51, 54 add 20 to above
n=12, 13, 14,15,16, 17, 18 divide above by 3
the youngest sensible ages would be smith 34 sawyer 54 and rein 18